OpenStudy (anonymous):

Which is a solution to the equation 5x + 2y = –1? A. (–2, 5) B. (1, –3) C. (–1, 4) D. (2, –4)

OpenStudy (anonymous):

@NeedHelpQuick546

OpenStudy (anonymous):

it's B

OpenStudy (anonymous):

Thank You!

OpenStudy (anonymous):

i can show u how to do it

OpenStudy (anonymous):

sure why not

OpenStudy (anonymous):

5*1 = 5 -3*2 = -6 add them : 5 + -6 = -1 :)

OpenStudy (sweetburger):

alright so this is your notation (x,y) so lets say you have a equation this is y=2x+1 and you have an x and y value of (2,5) this is how you would plug them in 5=2(2)+1 so this works out to 5=5

OpenStudy (anonymous):

Thank you